OpenStudy (anonymous):

x/(4/5)d = x5/4d right? I can't remember the process though, can someone please explain it. Thank you.

OpenStudy (anonymous):

\[\frac{x}{(4/5)d} = \frac{x5}{4d}\]

OpenStudy (lgbasallote):

\[\huge \frac{x}{\frac{4}{5}d} \implies \frac{x}{\frac{4d}{5}} \implies x \times \frac{5}{4d} \implies \frac{5x}{4d}\]

OpenStudy (anonymous):

Ahhh Thanks lg, that's awesome. Thanks too, Nancy!

OpenStudy (lgbasallote):

<tips hat>